### v1

# Funding AI for Good: A Call for Meaningful Engagement

Abstract.
Artificial Intelligence for Social Good (AI4SG) is a growing area exploring AI’s potential to address social issues like public health. Yet prior work has shown limited evidence of its tangible benefits for intended communities, and projects frequently face inadequate community engagement and sustainability challenges. Funding agendas play a crucial role in framing AI4SG initiatives and shaping their approaches. Through a qualitative analysis of 35 funding documents—representing about \$410 million USD in total investments, we reveal dissonances between AI4SG’s stated intentions for positive social impact and the techno-centric approaches that some funding agendas promoted. Drawing on our findings, we offer recommendations for funders to scaffold approaches that balance both contextual understanding and technical capacities in future funding call designs. We call for greater engagement between AI4SG funders and the HCI community to support community engagement work in the funding program design process.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: XXXXX
1. Introduction
Figure 1. AI4SG aims to bring about positive, long-lasting social impact. Prior literature emphasizes that impactful innovations are the result of a balance of technology capacities and contextual understanding (right panel). A bias towards technology capacities leads to innovations without a real-world positive impact (left panel). We use this spectrum as a lens to investigate whether current AI4SG funding agendas orient towards more techno-centric or balanced approaches.
Artificial Intelligence (AI) for social good (AI4SG) initiatives seek to bring about positive and long-lasting impact in social domains such as public health, by leveraging data-driven technologies like machine learning (Shi et al., 2020; Li et al., 2021; Perrault et al., 2020; Rolnick et al., 2022; Tomašev et al., 2020). However, a growing body of evidence suggests that AI4SG initiatives do not necessarily yield the intended positive outcomes in real-world contexts (Vinuesa et al., 2020; Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Tate Ryan-Mosley, 2023). Research in Computer-Supported Cooperative Work (CSCW), Information and Communication Technologies and Development (ICTD), and Human-Computer Interaction for Development (HCI4D) provides insights into the socio-technical mechanisms underpinning challenges in delivering the tangible positive impact of AI. While existing work has largely focused on the roles and perspectives of project teams (often consisting of AI technologists and mission-driven organizations), emerging works point to the significant influence of funders in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023).
Funders are organizations that provide monetary support and facilitate interdisciplinary programs, such as private foundations, nonprofit organizations, and technology companies. Funders function not only as financial sponsors but also as agenda-setters who significantly shape the ecosystems of AI development (Widder et al., 2024; Gardner et al., 2022). In AI4SG specifically,  Lin et al. (2024) draw on community organizations’ perspectives and reveal funders’ prevalent influence in how AI4SG partnerships are formed, which problem spaces are prioritized, and what kinds of solutions are pursued during the early stages of projects. Prior work also points out that funding often incentivizes mission-driven organizations’ tendency to adopt data-driven innovations, even at the cost of their original social missions (Bopp et al., 2017; Lin et al., 2024). For AI developers and practitioners, short funding timelines often constrain their considerations of ethics and unintended consequences (Do et al., 2023; Moitra et al., 2022; Okolo and Lin, 2024). While prior work has established funders’ prevalent influence through the perspectives of mission-driven organizations and AI developers in AI4SG, we still lack a direct understanding of how funders frame AI’s role in social good problems and how they envision AI’s social impact might be achieved.
To fill this gap, we take up Laura Nader’s (1972) call to “study up” AI4SG by analyzing its funding agendas (Nader, 1972). By focusing on funding, our work seeks to draw attention away from assigning blame to AI4SG project teams, and towards recognizing the broader power structures they are embedded in. Since AI4SG builds upon prior “tech for good” efforts (Aula and Bowles, 2023), we specifically draw on lessons learned from CSCW, ICTD, and HCI4D. This prior literature has extensively examined the approaches that influence the outcomes of socially driven technology projects. It emphasizes that in order for innovations to be impactful in real-world contexts, they must be grounded in both a thorough understanding of the problem and access to appropriate technological capabilities (Figure 1). A spectrum from techno-centric to balanced approaches emerges from our synthesis of this rich body of work, which points to a strong association between where projects fall on this spectrum and project outcomes. A techno-centric approach assumes novel technology alone can solve complex social issues, often leading to suboptimal results and strained relationships between researchers and communities (e.g., (Ames, 2019; Karusala and Kumar, 2017; Walsham, 2017; Dada, 2006; Brown and Mickelson, 2019)). In contrast, a more balanced approach is grounded in a solid understanding of the community’s challenges and contexts through community-collaborative approaches (Cooper and Zafiroglu, 2024), and tends to result in more community buy-in and sustainable efforts (e.g., (Toyama, 2015; Gandhi et al., 2007; Burrell and Toyama, 2009)). Through the lens of this spectrum, our study investigates whether current AI4SG funding agendas orient towards a more techno-centric or a more balanced approach. Specifically, we ask the following research questions:
  • RQ1: How do funders frame social good, and what is AI’s envisioned role in advancing social good, as articulated in funding documents?
  • RQ2: What are funders’ expected outcomes of AI4SG projects?
  • RQ3: How do funders envision these goals might be achieved?
To address these questions, we conducted a qualitative analysis of AI4SG funding calls for proposals and grant announcements (hereafter referred to as funding documents). Our dataset encompasses 35 funding documents—19 funding calls and 16 grant announcements—representing about \$410 million USD in total investments (see Appendix A for a complete list). These programs span a range of efforts, broadly framed under the banners of AI or data science for “social good,” “development,” “climate change,” and other social domains. Our dataset includes funding programs from 23 distinct funders, including private technology companies, philanthropic foundations, nonprofit organizations, and government agencies based in English-speaking countries such as the United States.
Funding documents in our corpus overwhelmingly mobilized AI against large-scale social problems that cannot be fully addressed with any single intervention (notably public health and development). Yet several funding documents framed AI’s potential as deterministically beneficial and positioned AI interventions as solutions to these social problems while skirting contextual and stakeholder specificity. Some others scoped interventions to more well-defined sub-problems and explicitly acknowledged both AI’s benefits and risks. Expected outcomes of funding programs in our dataset frequently centered on technical deliverables, while a subset adopted broader success metrics attentive to benefits to the communities involved. Furthermore, although interdisciplinary collaboration was widely encouraged, funding documents varied in their emphasis and scaffolding of community engagement and who were considered “experts.” With a few exceptions, expectations for community engagement stopped at a consultation mode towards later stages of projects (e.g., implementation stage), where community stakeholders were expected to provide inputs but not able to shape the broader project scopes.
To the best of our knowledge, this study presents the first qualitative analysis of how AI4SG funding documents frame social good, envision AI’s role in advancing social good, and anticipate desired outcomes to be achieved. While prior work has pointed out AI4SG’s tendency towards more techno-centric approaches, we provide detailed insights into how both techno-centric and balanced approaches are instantiated in different components of funding agendas. Contributing to existing literature on “tech for good” initiatives, we first highlight the disconnects between the positive social impact that AI4SG aims for and the techno-centric approaches that some funding documents promoted (Section 5.1). We then build on the more balanced approaches and existing gaps we identified in our dataset and propose additional opportunities for designing future funding calls, including supporting relationship-building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, emphasizing the role that the HCI community can and should play in helping shape AI4SG funding agendas and design processes (Section 5.3).
2. Related Work
Prior HCI research on AI4SG has highlighted both the movement’s intention for positive social impact and its challenges (Section 2.1). Scholars have called on AI4SG practitioners to draw on lessons from previous “tech for good” projects, which emphasize that impactful innovations require a balance of appropriate technical capacity and deep contextual understanding (Section 2.2). While prior work points out AI4SG’s bias towards emphasizing technical capacity, our work contributes an analysis of how this bias is instantiated across different components of funding documents, while offering concrete entry points for deeper contextual understanding. Despite the central role that funders play in shaping the AI4SG ecosystem, there is limited scholarship examining how funders frame AI4SG and what approaches are envisioned by funders (Section 2.3). In this study, we fill this gap by examining AI4SG funding documents through the lens of existing literature on “tech for good” projects.
2.1. The Promises and Challenges of AI for Social Good
Artificial Intelligence for Social Good (AI4SG) has emerged as a growing field of research and practice focused on applying AI to complex social issues (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). Its stated goal is to deliver transformative, positive, and lasting impact across many social aspects (Perera, 2024; Cowls, 2021; Tudor et al., 2024). Some notable examples include projects in agriculture that predict crop yields using remote sensing data (You et al., 2017) and provide price forecasting for farmers in India (Ma et al., 2019). There is a growing interest in both use-inspired AI4SG research (Shi et al., 2020; IJCAI, 2024) and real-world implementations (e.g., (Data Science for Social Good Fellowship, 2024; ITU, 2024)).
While enthusiasm for AI4SG remains high, growing evidence shows that many projects fall short of delivering lasting positive social impact and often end at the publication stage without deployment (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). In high-stakes contexts such as public service screening or criminal justice, AI may be inappropriate altogether (Eubanks, 2018; SALLY HO and GARANCE BURKE, 2023; Pruss, 2023). Additionally, AI4SG is often shaped by promotional hype (Joyce and Cruz, 2024), magical thinking (Lupetti and Murray-Rust, 2024), and “enchanted determinism” (Campolo and Crawford, 2020), where the fascination with the mysterious capabilities of new technologies captures public attention toward techno-solutionism—where complex, structural challenges are reduced to problems amenable to AI interventions (Morozov, 2013). This could be exacerbated by AI’s transparency issues (especially in the Global South context that AI4SG projects are intended for (Okolo and Lin, 2024)). For instance, studies in India show a strong belief in “AI authority” over public institutions and lowered expectations for accountability (Kapania et al., 2022; Ramesh et al., 2022).
HCI scholars have started to examine the mechanisms behind AI4SG initiatives. We highlight two key insights below. First, the aspiration to employ AI for social good is not a recent phenomenon; AI4SG should draw lessons learned from prior “tech for good” movements (Aula and Bowles, 2023; Lin et al., 2024). ICTD and HCI4D have long explored the development and use of emerging technologies in marginalized communities, especially in the Global South (Walsham, 2017; Dell and Kumar, 2016; Avgerou, 2010). Second, while past work often focuses on investigating perspectives of project teams—typically AI developers and nonprofit organizations, emerging research highlights funders’ role in shaping interdisciplinary partnerships (Lin et al., 2024; Saha et al., 2022a; Erete et al., 2023). Through interviews with staff of community organizations participating in AI4SG,  Lin et al. (2024) found that funding agenda influences program goals, proposed solutions, and definitions of success. Focusing on nonprofit organizations, other studies show how funding agendas shape their incentives to adopt emerging tech, and collaboration choices (Bopp et al., 2017; Erete et al., 2016; Goldenfein and Mann, 2023). AI developers and researchers’ approaches to AI ethics and considerations of unintended consequences are also shown to be impacted by funding timeline (Do et al., 2023; Widder et al., 2023; Moitra et al., 2022). Emphasizing the need to examine how power structures impact technology development in social contexts, scholars have taken up Laura Nader’s (1972) call to “study up” by focusing on those who shape social systems and exert normative influence (Nader, 1972). Emerging scholarship in algorithmic fairness (Barabas et al., 2020), public sector AI (Kawakami et al., 2024a), and machine learning data (Miceli et al., 2022) demonstrates that examining those in power can reveal critical leverage points for more accountable AI development.
While prior work has highlighted funders’ significant influence—primarily through the perspectives of mission-driven organizations and AI developers—we still lack a direct understanding of how funders frame AI’s role in addressing social good challenges. In this study, we “study up” AI4SG funding documents, through which funders communicate their intentions and envisioned approaches to AI4SG.
2.2. Lessons from Prior “Tech for Good” Projects: Techno-Centric vs Balanced Approaches
Prior work in CSCW, ICTD, and HCI4D has established a substantial body of work on technology projects aimed at social impact. Despite existing areas of debate within this body of work, there is a broad consensus that impactful innovations require not only engineering expertise but also situated knowledge and community engagement. Review studies on ICTD projects specifically point out that a bias towards technical problems while overlooking community contexts often leads to limited adoption and community impact (Brown and Mickelson, 2019; Heeks, 2003; Dodson et al., 2012), whereas sustained community engagement throughout the project life-cycle is a key factor for success (Brown and Mickelson, 2019).
In this section, we synthesize this rich body of work and categorize characteristics of projects into a more techno-centric approach (bias towards technical solutions) and a more balanced approach (emphasizing both technical capacities and contextual understanding). We particularly focus on aspects of motivations, collaborations, longevity and process, and evaluations, because these aspects are closely influenced by funding agendas (Lin et al., 2024). While prior work has pointed out a techno-centric orientation of some AI4SG initiatives, this paper offers a granular analysis of how this orientation is instantiated across specific components of funding documents.
2.2.1. Techno-Centric Approaches
A techno-centric approach, as defined by  Dodson et al. (2012), “place[s] the technology at the center of the intervention by giving prominence to technical features and capabilities.” It is rooted in “techno-solutionism,” where complex, structural challenges are reduced to problems amenable to technical interventions (Morozov, 2013), often under the false assumption that these solutions are neutral or apolitical (Feenberg, 2002; Winner, 2010). While leveraging emerging technology as a silver bullet to social problems appears optimistic and well-intentioned, Berlant’s concept of “cruel optimism” points to the cruel side of that optimism: promising social transformation while deepening surveillance, reproducing bias, or ignoring community needs (Berlant, 2011). These initiatives often fail due to “design-reality gaps” (Heeks, 2002): a mismatch between system designers’ perceptions of local contexts and local actuality.
In “tech for good”, techno-centric projects are driven by an “innovate at all costs” mentality, leading to the creation of technologies with little maintenance support and local capacity building afterward (Saha et al., 2022b). It incentivizes a top-down design approach, where the solutions are typically designed by external technologists without adequate engagement with the local communities (Walsham, 2017; Dada, 2006; Harris, 2016; Dodson et al., 2012), especially during the decision-making stages of projects (Saha et al., 2022a). It could also lead to the phenomenon of “bungee jumping research” (Dearden and Tucker, 2016), characterized by brief, superficial engagement (Saha et al., 2022b; Sultana et al., 2019; Dada, 2006). Development of the solutions tends to follow a linear and short-term process: building the technology, deploying it, and moving on to the next project, where researchers often leave after their work is published, making it difficult to sustain the systems they have introduced (Taylor et al., 2013; Dada, 2006; Walsham, 2017). Last but not least, evaluations of techno-centric projects constrain assessments of success to technical dimensions, narrowly focusing metrics such as model accuracy or the production of tangible artifacts like software (Dodson et al., 2012; Lin et al., 2024).
2.2.2. Balanced Approaches
A balanced approach is grounded in the understanding of problems faced by specific communities, in addition to leveraging emerging technology’s capacities. Community-collaborative approaches (CCA) are central to a balanced approach. A systematic review by Cooper et al. defines CCA as projects that “involve collaboration with community stakeholders as co-researchers throughout the research process while investigating complex problems of concern to the community (e.g., social and health-related), and developing novel technologies or sociotechnical solutions” (Cooper et al., 2022). CCA includes work in participatory design (e.g., (Delgado et al., 2023; Bondi et al., 2021)), community-based participatory research (e.g. (Wan et al., 2023; Liang et al., 2023; Strohmayer et al., 2017)), and action research (e.g., (Hayes, 2011; Le Dantec, 2016; Lemaire and Muñiz, 2011; Balestrini et al., 2014)). Projects that follow CCA are more likely to create lasting social impact because they have community support and input from the start (Gandhi et al., 2007) and amplify positive values already present in the community (Toyama, 2015)Brown and Mickelson (2019) emphasize the necessity of “the full inclusion of all relevant parties in every aspect of the project” to ensure project success.
A balanced approach emphasizes that community members’ lived experiences are a legitimate form of knowledge within their specific contexts (Erete et al., 2023; Cooper et al., 2022), on par with technical expertise. This emphasis is grounded in the concept of epistemic justice (Fricker, 2007), which calls for fairness in recognizing who is seen as a knower and what forms of knowledge are valued. In “tech for good” projects, this means embracing modalities of knowledge sharing that go beyond traditional academic outputs, such as storytelling and community observations (Erete et al., 2023; Harrington et al., 2019a; Zegura et al., 2018)Zegura et al. (2018) demonstrate that valuing the community’s local knowledge is not only crucial for the quality of data and the subsequent data analysis, but also generates “care” among researchers and community collaborators (Zegura et al., 2018). Relationships among collaborative teams are critical for the feasibility and sustainability of community-based projects (Dell and Kumar, 2016; Kumar and Dell, 2018). Sustainable relationships with communities can also be cultivated through regular presence, trust-building, and iterative design processes, and they enable project agendas to emerge organically from (rather than being imposed on) communities (Dearden and Tucker, 2016; Carroll and Rosson, 2013). Lastly, a balanced approach extends evaluation beyond technical outputs to include community outcomes such as increased technical capacity (Zegura et al., 2018), impact on everyday practice (Taylor et al., 2016), and “empowerment of the participants and the knowledge gained by the technical designers” (Drain et al., 2021).
Prior work has started to evaluate the extent to which an AI project grants agency to community stakeholders in different stages of projects (Corbett et al., 2023; Feffer et al., 2023; Delgado et al., 2023). For example, Delgado et al. (2023) develop the Parameters of Participation framework that lays out the dimensions of participatory AI work, categorizing projects into a spectrum of four modes—consult, include, collaborate, and own—and three dimensions—participation goal, scope, and form. The authors then use the framework to analyze 80 research articles, revealing that existing participatory AI projects focus more on community consultation and less on collaboration and stakeholder ownership (Delgado et al., 2023).
Prior work has also highlighted some limitations of CCA (Pine et al., 2020; Erete et al., 2023; Harrington et al., 2019b). For example, nonconsensual, short-term, and non-contextual-specific participation would lead to the exploitation of impacted communities through “participation washing” (Sloane et al., 2020). Sloane et al. argue for recognition of participation as work that is valuable for machine learning development, including data work and context-specific consultation (Sloane et al., 2020). In AI4SG specifically, Lin et al. document the contributions that community organization staff make to AI4SG projects, including data work and participating in regular feedback meetings (Lin et al., 2024). However, this labor is often invisible or overlooked by downstream AI developers and the public (Crain et al., 2016; Gray and Suri, 2019; Crawford, 2021; Sambasivan et al., 2021). A recent body of work has started to quantify invisible labor in crowd work (e.g., managing payments and hyper-vigilance) (Toxtli et al., 2021) and online volunteer moderators (e.g., approving content and managing users) (Li et al., 2022), revealing that such unseen tasks make up a substantial portion of the overall work.
2.3. Gaps in Our Understanding of AI4SG Funders and Funding Agendas
Despite the central role that funders play in shaping the AI4SG ecosystem, there is limited scholarship examining how funders frame AI4SG and what approaches are envisioned by funders. An emerging body of work has started to understand the perspectives, incentives, and challenges of grantmakers in the broader “tech for good” space. For example, through interviews with eight decision-makers in development projects, Saha et al. (2022a) found that community participation is rarely a part of the grant decision-making process, and it is not a requirement for applicants either. Some barriers to incorporating community voices include limited access to local organizations and communities and overconfidence in their knowledge about the local contexts (Saha et al., 2022a). Relatedly, Gardner et al. (2022) point out that a lack of internal expertise in AI and AI ethics creates barriers for funders to incorporate Trustworthy AI principles. Studies on other influential decision-makers, such as policymakers in government agencies, point out factors influencing their decision-making process around innovative technologies, including learning from neighboring agencies’ experiences, reliance on national surveys (which omit marginalized communities’ views of their needs) (Saha et al., 2022b). These decision-makers also face obstacles in effective community engagement, including power and knowledge differentials and inadequate infrastructural support (Kawakami et al., 2024a).
Focusing on private tech companies specifically, some recent works have started to critically examine investment in AI4SG (Nost and Colven, 2022; Lukacz, 2024; Henriksen and Richey, 2022). For example, Nost and Colven (2022) view Microsoft’s AI for Earth program as a form of greenwashing and “philanthrocapitalism”, where the company uses the initiative to enhance its reputation as a socially responsible corporation, while simultaneously contributing to environmental harm through its broader business operations, including its reliance on fossil fuels for data centers. Lukacz (2024) further argues that Microsoft’s narrative of democratizing AI and environmental data obscures the power they wield in shaping the technology and its impacts. Henriksen and Richey (2022) examine Google’s philanthropic efforts, arguing that they serve dual purposes: advancing social causes while simultaneously expanding Google’s market influence and integrating its technologies into more areas of life. Comparing “Big Tech” companies with the “Big Tobacco” industry, Abdalla and Abdalla (2021) examined “Big Tech” funding’s influence on the broader academic landscape (though not specific to AI4SG), impacting events and research questions that “put forward a socially responsible public image” of tech companies.
3. Methods
To examine how funders conceptualize AI4SG and what approaches they envision in AI4SG, we analyzed publicly available funding documents. In this section, we detail our approach for identifying and analyzing AI4SG funding documents (Subsections 3.1 and 3.2) as well as our team’s positionality (Subsections 3.3). We discuss limitations to our approach and opportunities for future work in Subsection 5.4.
3.1. Qualitative Dataset Development
Definitions
In this paper, we define “funders” as organizations that provide monetary support to AI4SG, such as philanthropic foundations, nonprofit organizations, government agencies, and private technology companies. While individuals may also fund AI4SG projects, we focused on organizations because they are more likely to provide public articulations of their funding agendas. We define “funding programs” as specific AI4SG initiatives, such as a research grant on the intersection of climate change and AI. We define “funding documents” as funding calls for proposals and grant announcements through which funders communicate their intentions and envisioned approaches in AI4SG. While funding calls typically outline more detailed guidelines, grant announcements present a more concise overview of an initiative’s aims, scope, and thematic priorities. We consider both as funding documents because they offer complementary insights into funders’ intentions.
Sampling Approach
We used the Google search engine to identify publicly available AI4SG funding documents published prior to August 2025. Following the approach in Wong et al. (2023) when searching for AI ethics toolkits, we sought to emulate how AI4SG stakeholders (e.g., research teams, community-based organizations, individual practitioners) might seek funding opportunities. This approach also led us to discover and examine webpages from grant databases, such as nsf.gov/funding/opportunities.
We conducted the search with keywords indicating 1) the emphasis on social good (“social good,” “social impact,” “climate change,” “public health,” “justice,” “equity,” and “community”), 2) the methodological orientation (“Artificial Intelligence,” “AI,” “Machine Learning,” “ML,” or “Data Science”), and 3) the focus on funding (“funding,” “fund,” “award,” or “grant”). The query returned 296 links to web pages. The links led to calls for proposals, announcements about the grants, blog posts about AI4SG projects, academic publications in AI4SG, academic departments and labs working on AI4SG, and so on. Using snowball sampling (Naderifar et al., 2017), we followed references within the initial links to identify additional funding programs. For example, a university funding opportunities page [^1] https://e-hail.umich.edu/funding-opportunities/ accessed Aug 2, 2025 listed multiple AI4SG-related programs not captured in the original search. We additionally examined these funding programs for inclusion.
Inclusion and exclusion criteria
We identified funding documents for inclusion using criteria reflecting three dimensions of our search terms: the emphasis on social good, the methodological orientation, and the focus on funding. The following three criteria were developed through iterative discussions between the first two authors, after jointly reviewing a random subset of 50 links from the search results.
  • Funding programs must focus on addressing societal challenges such as humanitarian resource allocations, public health, and climate change. This excluded funding programs that focus on using AI for natural science discoveries (e.g., astronomical sciences (Simons Foundation, 2023), chemistry (NWO, 2024)) or AI foundational and theoretical research (NSF, 2024).
  • Funding documents must specify an orientation towards developing, using, or evaluating AI, Machine Learning, Data Science, or other data-driven technologies for social applications. This excluded projects on AI education (e.g. (NSF, 2023)), which focus on educating people about AI instead of exploring AI as an application in the education field (e.g., AI assistants in classrooms).
  • Funding documents must be published directly by the funders (rather than grantees), and contain details about the funding programs, including monetary funding amounts, durations of funding support, and eligibility criteria. This excluded initiatives that solely offer technical training, computing resources, or visibility without direct monetary support, such as accelerators (e.g. (APRU, 2025)), competitions (e.g. (DrivenData, 2025)), and recognition awards (e.g. (Maheshwari, 2025)). Furthermore, given our interest in examining eligibility criteria across a broad spectrum of targeted applicants, we also excluded funding programs restricted to internal applicants, such as a university research seed grant available only to its faculty and students (e.g. (Penn State University, 2025)).
Using these criteria, the first author then examined the remaining links, and the second author reviewed the selections. The first two authors then resolved any disagreements. We narrowed down to 35 funding documents (19 funding calls and 16 grant announcements) for in-depth qualitative analysis. The funding programs represent 23 funders and around \$410 million in investments in aggregate. We also included supplementary materials referenced in funding documents in our qualitative dataset, as they offer important details about the funding programs, such as Frequently Asked Questions and application forms. For example, we supplemented the Google AI Impact Challenge funding call (Google, 2018) with its application guide (Google, 2019), which laid out selection criteria and application questions.
Dataset description
Table LABEL:tab:funding_doc_mr in Appendix A provides the list of funding documents and information about the funding programs (e.g., funding amount, duration, supporting funders, target regions). Our dataset covers a wide range of programs from 2017 to 2025, aiming at both real-world implementation and use-inspired research. These funding documents were provided by government agencies (11 funding documents), philanthropic foundations (10), nonprofit organizations (7), and private technology companies (8) based in English-speaking countries. These funders were headquartered in the United States (26), the United Kingdom (5), Canada (4), Sweden (3), Germany (1), Ireland (1), and Australia (1). These funding programs targeted applications from countries including the United States (5), the United Kingdom (1), Ireland (1), and South Africa (1), as well as broader regions including globally (14), Africa (9), Latin America (2), and Low- and Medium-Income Countries (2).
3.2. Qualitative Analysis
We conducted a thematic analysis (Braun and Clarke, 2006, 2019) of the funding documents through an iterative coding process. In the first stage of coding, the first two authors coded basic information about the funding programs, including the methods, the social issues, the duration of projects, the intended applicants, and the amount of funding. In the next coding stage, the first two authors independently conducted an open-ended inductive qualitative analysis of three randomly chosen funding documents and discussed the coding strategies. Through iterative discussions, the two authors identified some initial themes, which helped refine our research questions and narrow the scope of analysis to four main components of funding documents: 1) Background, 2) Expected Outcomes, 3) Eligibility Criteria and Team Composition, and 4) Funders’ Support. We also kept detailed memos as we coded the themes to document our observations. In the third coding stage, the first author coded all the funding documents, with our research questions in mind. Each funding document’s codes were second reviewed by another author in our team. The first author then resolved any discrepancies with the rest of the team through online discussions. All authors then conducted iterative affinity diagramming on an online whiteboard and discussed emerging high-level themes from the codes. We collaboratively characterized these high-level themes into a spectrum between a more techno-centric approach or a more balanced approach using the lens of prior literature synthesized in Section 2.2. Appendix B provides screenshots of our coding process and affinity diagrams.
3.3. Positionality
The authors’ collective prior research and practice in public sector AI and AI for social good have led us to believe that funding agendas are important to study. We started with the basis that power imbalances exist within AI4SG partnerships and sought to examine power from the very beginning of AI4SG projects: when they were conceptualized for funding. While the authors are from diverse ethnic and cultural backgrounds, all are researchers in academia living in the United States and conduct research in English.
4. Findings
In what follows, we report on themes in different components of funding documents and highlight rhetoric leaning towards a more techno-centric or balanced approach to AI4SG. We begin with the Background (Section 4.1), where we discuss how funders frame “social good”, “AI”, and how AI is envisioned to help address social good problems (RQ1). In the Expected Outcomes and Evaluations (Section 4.2), we discuss funders’ envisioned outcomes, often communicated through evaluation criteria (RQ2). In the Eligibility (Section 4.3) and Support from Funders (Section 4.4) sections, we shed light on how funders envision “AI4SG” can be achieved and by whom (RQ3). Table LABEL:tab:funding_doc_mr in Appendix A provides a detailed overview of our dataset, including funding amount, duration, targeted regions, funder types, and target regions.
4.1. Background: Motivations and Beliefs Underlying AI4SG Funding Documents
Most AI4SG funding documents in our corpus focused on addressing large-scale, complex social issues, including public health (6, e.g. [C17, C19]), development (6, e.g. [A7, A8]), climate change (3, e.g. [C11, C12]), and mental health (2, [C8, C9]). These social problems have been categorized as “wicked problems”—complex, large-scale, and often ill-defined problems created by multiple, interdependent, evolving issues that any single intervention cannot fully address (Gurumurthy and Chami, 2019; Lönngren and van Poeck, 2021). However, some funding documents depicted AI-based interventions as a solution to these large-scale social problems. They emphasized AI’s transformative benefits deterministically while glossing over or omitting the specific contexts and stakeholders that AI could be beneficial for (Sections 4.1.1). Some funding documents, on the other hand, were grounded in a specific context within a large-scale social issue, sometimes providing a detailed literature review. When describing AI’s role in addressing the social problems, some funding documents acknowledged both the potential benefits and harms of AI in these specific contexts (Section 4.1.2).
Figure 2. Themes related to rhetoric of motivations for AI4SG that lie across the spectrum between a more techno-centric versus balanced approach. The positions of the themes do not represent their prevalence.
4.1.1. More techno-centric
Emphasizing AI’s transformative potential with limited contextual grounding. Funding documents in our corpus rarely provided an explicit definition of what “AI” is 222with the exception of C3 and C13 (both in the Frequently Asked Questions sections). Yet several funding documents emphasized AI’s role as a potential solution to complex (and often vaguely defined) social issues, without instantiating exactly how AI could be beneficial and for what contexts. For example, UNICEF Innovation Fund Call for Data Science & A.I explicitly looked for “data science, machine learning, artificial intelligence or similar technology solutions” that would broadly “benefit humanity” (C2). The call provided example questions to elaborate on what benefiting humanity with AI means, though the application area remains broad and vague while pre-determining the use of AI as a solution: “Are you using new sources of data such as satellite imagery or social media and applying data science or artificial intelligence techniques to understand the physical world that we live in?” Under this question, the call then further listed out examples of how AI might be beneficial, such as “[a]pplying deep learning to analyze satellite imagery and automatically map infrastructure such as schools, health centers, roads or cell towers.” While this example provided a more specific example of how AI could be used for a specific task (using satellite imagery to automatically map infrastructure), the call did not make clear why automatically mapping infrastructure like schools might be beneficial, and for whom it would be beneficial.
Similarly, in an official announcement of Microsoft’s AI for Humanitarian Action program (A14), AI was depicted as a “game changer, helping save more lives, alleviate suffering and restore human dignity by changing the way frontline relief organizations anticipate, predict and better target response efforts” in humanitarian crises, as “[frontline relief organizations’] work by definition is often reactive and difficult to scale.” In this depiction, AI played a transformative role in helping relief organizations “anticipate, predict and better target” efforts. However, this depiction omitted what specific challenges relief organizations faced, and which AI could (or could not) address.
Promoting urgent dissemination of AI to resource-constrained contexts. Several AI4SG funding programs were designed to diffuse the benefits of AI technologies to developing countries and the nonprofit sectors from Western countries and the private technology sector. This diffusion of AI was seen to be “urgent,” so that under-resourced communities in developing countries or the nonprofit sectors could keep up with the rapid development of AI. Some funding programs from international development agencies in Western countries, such as the AI for Development Africa program funded by IDRC (Canada) and SIDA (Sweden), emphasized the importance of ensuring “developing countries are not left behind in the AI revolution” (A9). Some funding programs offered by private tech companies and philanthropic foundations similarly extend this narrative to the nonprofit sectors: “While private-sector businesses have been building and deploying artificial intelligence for years, most organizations in the nonprofit, civic, and public sectors have yet to robustly apply these techniques towards the complex challenges they address” (A6). As Patrick J. McGovern Foundation’s grant announcement (A1) stated, enabling nonprofits to use AI and data “have never been more urgent.” In these funding documents, the rhetoric of AI4SG led with the assumption that AI’s benefits could be directly transferred to non-Western countries and the nonprofit sectors.
4.1.2. More balanced
Situating AI’s potential within contextual understanding of scoped social problems. Some funding programs in our corpus were motivated by not only AI’s potential capacities, but also more concrete contextual needs (within a larger social domain). For example, the Call for Innovation Research Grants on Responsible Artificial Intelligence for Climate Action (C11) started with a broad focus on climate change in Africa. Within this broad focus, the call provided background information specific to the African context: “Drought is the most studied hazard across Africa. A large part of Africa is susceptible to drought, […]. Wildfires have been more localized particularly in the recent history in Zimbabwe and Lesotho.” The call then motivated proposals of AI use targeted towards “understanding the drivers of droughts in Africa” and predicting drought characteristics. In the application form, the funding call then provided opportunities for applicants to specify who the proposed projects would benefit and how: “[w]ho are the target beneficiaries and what process has been undertaken to engage them in the development of the proposed action? What are the needs and challenges of each of the target groups identified and how do these relate to the action proposed?” (C11)
Another notable example is the Wellcome Mental Health Data Prize: UK and South Africa (C8), which solicited data projects that addressed a more specific problem space within mental health: “anxiety and depression in young people.” The funding call had a regional focus on the UK and South Africa, and the funder worked with data holders from these regions to provide exemplar datasets to applicants. In addition to exemplar datasets, the funding call built on existing “evidence for different active ingredients deemed to help prevent, treat, and manage anxiety and/or depression in 14 to 24-year-olds globally.” Applicants were then both encouraged to build on these existing findings or explore “research that expands the scope of the existing research.” In both cases, the funding calls were motivated by not only the potential of AI, but also more specifically scoped problems within a larger social domain.
Framing AI as one tool among many and acknowledging its potential benefits as well as risks. A few funding calls provided a more nuanced depiction of AI capabilities, acknowledging both its potential benefits and risks, as one part of broader problem-solving efforts. For example, the Mozilla Foundation’s AI and Environmental Justice Awards (C7) began with the statement that “AI’s relationship with our current climate emergency is complex. The technology produces a vast carbon footprint, and can be used to hasten the extraction and exploitation of natural resources. Yet AI also has the potential to drive environmental justice at scale.” The funding call (C7) also posed questions (rather than assumptions) about whether AI may be helpful for specific use cases. Specifically, the funding call asked: “Could AI technologies be used to catalog and analyze environmental degradation, land and water use, pollution and other environmental impacts?” It explicitly explored AI as “a part of the solutions to these issues,” rather than assuming that AI alone could be a solution.
4.2. Expected Outcomes and Evaluations
Funding programs in our dataset commonly aimed to produce “innovative,” “scalable,” ”feasible,”, and “sustainable” AI-based interventions or “novel” research outputs, often communicated through proposal evaluation criteria. Some funding documents emphasized the technical dimensions of desired outcomes, and some sought to showcase the benefits of AI through research outputs (Section 4.2.1). In contrast, some funding documents evaluated desired outcomes beyond the technical to emphasize benefits to the community involved. Some documents also encouraged flexibility in pursuing research novelty and evaluating the negative impacts of AI technologies (Section 4.2.2).
Figure 3. Themes related to funders’ expected deliverables for AI4SG projects that lie across the spectrum between a more techno-centric versus balanced approach. The positions of the themes do not represent their prevalence.
4.2.1. More techno-centric
Grounding evaluation in technical dimensions over community outcomes. Many funding programs in our corpus aimed at deploying new or scaling existing AI systems to tackle social challenges. While funding documents rarely specified success metrics for deployed systems and who got to define success, expected outcomes were often implied through the proposal evaluation criteria. Some funding documents underscored technical dimensions of success while placing less emphasis on social processes, community empowerment, or project sustainability beyond the technology itself. For example, the Google AI Impact Challenge application guide (C3) specified criteria of successful applications, including “Impact,” “Use of AI,” “Feasibility,” and “Scalability.” “Impact” asked for “a clear plan to deploy the AI model for real-world impact,” framing societal benefit largely in relation to the effective application of AI. “Use of AI” explicitly required applicants to integrate AI technology. While “Feasibility” acknowledged the importance of partners and domain experts for the successful implementation, the criterion placed particular weight on the technical features of the proposed plan, asking for “a plan to access a meaningful dataset and technical expertise to apply AI to the problem.” The “Scalability” criterion valued projects that could expand the growth and replication of AI systems, suggesting the use of “metrics around speed, accuracy, cost, or scalability” in an application question.
“Demonstrating” the benefits of AI with research outputs. Other funding programs in our dataset aimed at advancing use-inspired AI research (without a requirement for AI deployment), including Amazon’s Data for Social Sustainability grant (C6) and Microsoft’s AI for Earth program (A12 and A13), both of which came from the research arms of private tech companies. Most funding documents expected the dissemination of research outputs and contributions to the research community, often in the form of “publications, presentations, code and data releases, blogs/social media posts, and other speaking engagements” (C6), while placing limited to no emphasis on disseminating research outputs with contributors or partners in the community. Moreover, some funding programs sought to demonstrate rather than evaluate the uses of new AI technologies. For instance, the press announcement for the AI for Earth program (A12) described that they intended to “demonstrate how AI can deliver results more rapidly, accurately, and efficiently” and “encourage others to innovate based on the power and potential of AI.” This statement was grounded in the assumption that AI would be deterministically beneficial.
4.2.2. More balanced
Grounding evaluation in benefits to the local community beyond the technical aspects. A few funding programs in our corpus went beyond emphasizing technical deliverables and explicitly highlighted benefits to the community as part of their evaluation framework. For example, Lacuna Fund’s call for proposals related to datasets for Natural Language Processing (C14) and Antimicrobial Resistance (C15) specified several evaluation criteria in addition to technical quality and feasibility of proposals, particularly highlighting dimensions of “Participatory Approach” and “Sustainability & Communications.” The “Participatory Approach” criterion required applicants to “ensure sustained maintenance and usage of the dataset by the local community” and to “share project outputs and benefits with data providers and/or the community,” emphasizing the tangible benefits of the projects to the community involved. Similarly, the “Sustainability & Communications” criterion encouraged applicants to embed community-oriented structures for ongoing impact, for example, by “organizing a workshop on dataset sustainability with interested stakeholders; or establish a sustainability committee.”
Evaluating AI’s impact and encouraging flexible expectations of novelty in research. Some funding calls aimed at advancing AI research explicitly supported assessments on the direction of AI’s potential impact (rather than just the scale of its benefits). For example, the Wellcome Mental Health Data Prize: UK and South Africa (C8) encouraged applicants to consider the risks posed by data biases in AI technologies and requires applicants to address data quality, bias mitigation, and ethics in their proposals. Applicants must consider “risks or issues associated with the proposed timelines and plans for mitigation,” with this criterion accounting for 10% of applicant evaluation.
Some funding calls also expected research novelty more flexibly to ensure the exploration of AI-based interventions was suitable for the specific contexts. For example, the Wellcome Mental Health Data Prizes (C8 and C9) clarified that “it is not expected that the team must innovate in their methodology if this is not appropriate for their research.” Similarly, the Climate Change AI Innovation Grants (C13) specified that “it is perfectly acceptable to use a previously published ML algorithm or architecture,” and grounded the exploration of AI on a solid contextual understanding: “Projects should employ or address AI/ML in a way that is well-motivated and well-scoped for the problem setting, which may or may not include the development of novel techniques.”
4.3. Eligibility Criteria and Team Composition
Funding programs in our corpus commonly viewed interdisciplinary collaborations as valuable for achieving the aims of AI4SG. However, some funding calls had requirements that were biased towards (and could only be realistically met by) teams with existing AI capacities (Section 4.3.1). Funding calls also differed in how they scaffolded involvement of different stakeholders, as well as who was seen as a “domain expert” (e.g., those with scholarly knowledge versus lived experiences). Overall, we observed that many funding documents that encouraged involvement of community stakeholders centered around consultative modes of participation and during later AI project stages (i.e., implementation phase) (c.f. (Delgado et al., 2023)) (Section 4.3.1). In Section 4.3.2, we spotlight a few funding documents that went beyond the consultative mode and encourage community ownership.
Figure 4. Themes related to edibility criteria of AI4SG funding calls that lie across the spectrum between a more techno-centric versus balanced approach. The positions of the themes do not represent their prevalence.
4.3.1. More techno-centric
Targeting pre-existing technical capacity while giving less weight to contextual knowledge. Several funding calls in our corpus designated single-applicant recipients, making equal ownership among interdisciplinary collaborations challenging. While most single-recipient grants in our corpus still allowed for subcontracting and encouraged partnerships, a few funding programs’ eligibility criteria privileged organizations with pre-existing AI expertise, while omitting requirements for community engagement or the integration of contextual knowledge. For example, the UNICEF Innovation Fund Call for Data Science & A.I. (C2) looked explicitly for “for-profit technology start-ups” that had a working prototype with promising results. The intention of the grant was to “seed an exploration with a company that already exists, that has a strong team and a prototype, and that can be profitable, and successful in the space of data-science.” While the aim for social impact was clear in the funding call, it did not mention an expectation for applicants to engage with community or domain experts to ensure adequate contextual understanding. Some grant announcements also signaled a focus on funding organizations with existing AI capacities. For instance, the Rockefeller Foundation announced a \$5.5 million investment in “satellite data & AI to accelerate economic development and climate resilience in Africa” (A5). The fund was granted to two organizations consisting of researchers (based in the U.S.) with “extensive experience building […] datasets, predictive analytics models, and software platforms” (A5).
Community engagement limited to consultation in later stages of projects. Across many funding documents in our corpus, we observed that expectations for community engagement—in cases where they existed at all—seldom surpassed the consultation mode, where impacted communities and intended end beneficiaries could provide input, but they were not expected to shape the projects’ overall scope and purpose or play a central role across the project life cycle. For example, some funding calls only encouraged community consultation in the implementation stage, rather than the early stages of problem identification and solution development. For example, the Google AI Impact Challenge (C3) specifically asked applicants to “identif[y] the right partners and domain experts needed for implementation.” The funding call explained that involving these stakeholders during implementation could help “deploy the AI model [to support] real-world impact,” where expected community engagement primarily served the purpose of AI deployment.
4.3.2. More balanced
Requiring domain expertise or contextual knowledge in project teams. Some funding calls built interdisciplinary collaborations into team structure requirements, requiring that a variety of expertise was represented in the core project team. For example, the Artificial Intelligence (AI) for Societal Good Science Foundation Ireland Future Innovator Prize (C1) determined the structure of a core project team, comprising a “Team Lead,” a “Team Co-Lead,” and a “Societal Impact Champion.” The team lead and co-lead were expected to provide technical leadership while the “societal impact champion” was envisioned to “have appropriate experience in areas relevant to the societal impact focus of the application.” Similarly, the Artificial Intelligence (AI) innovations in Maternal, Sexual and Reproductive Health (MSRH) call (C10) mandated that teams include a domain expert who would ensure that gender equality and inclusion were appropriately understood and considered, specifically encouraging representation from “local gender equality or women’s rights organizations or underrepresented groups.”
Some funding calls specifically targeted applicants from a geographic region of interest, emphasizing locally-owned projects. The Catalyzing Equitable Artificial Intelligence (AI) Use call for proposals (C8), for instance, had a specific focus on the use of Large Language Models in low- and middle-income countries (LMICs) and required that proposals were “led by investigators in LMICs.” The funding call further clarified that: “priority will be given to proposals that demonstrate at least 80% of the funding is going to LMIC institutions […]”. This requirement specified that funding would directly flow to organizations from the regions of interest.
Broadening definitions of “domain experts” and encouraging deeper community engagement beyond consultation. A few funding calls highlighted the need to engage with community organizations or members from the affected communities directly. For example, the Wellcome Mental Health Data Prize: Africa (C9) application guide required that “individuals with lived experiences” of mental health problems were part of the multi-disciplinary team, alongside “data scientists and mental health researchers”. The funding call’s emphasis on “integrating meaningful engagement of community and implementing partners at every stage of the research” was built into its evaluation criteria, where engagement accounted for 10% of an application’s score. Additionally, the funding call emphasized that “contextual understanding” and experiences in community involvement were necessary skills alongside technical proficiency, asking applicants to highlight past experiences working with communities in the application forms.
A few funding calls required inclusion of community stakeholders as key decision-makers across the project pipeline (beyond consultation). For example, the “Societal Impact Champion” in project teams of the Artificial Intelligence (AI) for Societal Good Science Foundation Ireland Future Innovator Prize (C1) was expected to “identify and validate challenges in addition to advising on solution development,” making sure that the projects were grounded in real-world needs. Beyond problem identification, the funding call also emphasized that “key measures of success [are] identified through engagement with stakeholders and beneficiaries.”
4.4. Funders’ Support and Connections
Funding calls in our dataset supported projects ranging from three months to five years (with a median of one and a half years). The longer-term programs were typically funded by governmental agencies (e.g., C17 and C18 by NIH, and C19 by NSF). Overall, we observed a lack of provisions for post-deployment funding support. Many funding programs readily provided non-monetary support that was biased towards technical aspects of the projects (Section 4.4.1). Some other funding programs provided broader forms of support beyond the technical, sometimes through partnerships with regional organizations (Section 4.4.2). However, across the funding documents in our corpus, there was a notable lack of guidance and training for applicants to develop skills in community and stakeholder engagement.
Figure 5. Themes related to the support that AI4SG funders offered that lie across the spectrum between a more techno-centric versus balanced approach. The positions of the themes do not represent their prevalence.
4.4.1. More techno-centric
Lacking post-deployment funding support. While sustainability was a consideration for many funding calls aiming at AI deployment, there was a notable lack of provisions for post-deployment funding support. For example, the Google AI Impact Challenge (C3) asked applicants how they would plan to “sustain and grow the impact of this work beyond this grant” but did not clarify whether future funding would be available to support this post-deployment work. Some funding calls indicated that teams needed to sustain their deployed AI systems on their own, independent from the funders. For example, IDRC and SIDA’s AI for Climate Action Innovation call for proposals (C11) expected applicants to “source for funding for sustainability of the innovation” themselves.
Biasing towards technical support. Across the funding programs in our dataset, we also observed a notable lack of training and resources towards building contextual understanding, especially for community engagement, while technical training and access to computing resources appeared to be a standard offering. Private technology companies, in particular, often provided technical support using their own infrastructures and products. For example, Microsoft’s AI for Earth program offered “access to cloud and AI computing resources [and] technology trainings” (A12 and A13), later clarifying that this included its own “Azure compute time and our data science virtual machine offerings on Azure.”
4.4.2. More balanced
Tailoring support activities to the needs of applicants. Some other funding programs provided broader forms of support, beyond just technical resources and training. For example, as part of the Wellcome Mental Health Data Prize (C8 and C9), Wellcome Trust offered a web-based platform to facilitate knowledge-sharing across teams. The funder also facilitated matchmaking events where “individuals with different knowledge, skills and backgrounds will be introduced to collaborate and develop a proposal together” to enable interdisciplinary collaborations. In another funding call (C7), the Mozilla Foundation described a broad array of support that funded teams would receive: “While the specific types of support are tailored to the unique needs of the awardees and cohorts, providing opportunities for peer learning and community, mentorship, and communications support are key elements of our accompaniment approach.” However, none of the funding documents in our corpus explicitly mentioned training resources and support for local community engagement and participatory design.
Partnering with regional organizations where the projects are based. A few funding programs involved regional partners in the design of the funding calls (C8-C12). A notable example was the design process of Wellcome Trust’s Mental Health Data Prizes (C8 and C9). After the first round of the data prize in the UK and South Africa (C8), the funder reflected that “it did not attract as many applications as expected from South Africa” (Wellcome Trust, 2023b). After conducting feedback sections (Wellcome Trust, 2023a), the funder worked with a regional partner, the African Population and Health Research Center, to develop the funding call in Africa (C9). The regional partner played a “primary role” in “prize management” (e.g., developing thematic areas and leading the application selection process), “capacity building” (e.g., providing training in data and mental health research), “matchmaking” (e.g., reach communities and potential participants in Africa), and “landscape analysis of existing mental health research” (e.g., literature review on existing data sources) (C9). The partnership with the regional partner resulted in several key adjustments: providing a literature review on mental health issues in the African context, extending the overall timeline for project development, and indicating that a “lived experience expert” would be a member of the application selection panel.
5. Discussion
Our analysis presents a range of framing and approaches envisioned by AI4SG funders through funding documents. In this section, we first highlight the disconnects between the positive social impact that AI4SG aims for and the techno-centric approaches that some funding documents promoted (Section 5.1). Building on the more balanced approaches and existing gaps identified in our dataset, we then propose additional opportunities for designing future funding calls, including supporting relationship building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, emphasizing the role that the HCI community can and should play in helping shape AI4SG funding agendas (Section 5.3). We discuss the limitations of our work and lay out several future research directions in Section 5.4.
5.1. Revealing Dissonances Between Intended Outcomes and More Techno-centric Approaches
Many AI4SG funding programs in our dataset expected feasible and sustainable deployment of AI technologies or novel research advancement in understanding AI’s potential (Section 4.2). However, more techno-centric approaches promoted through some funding documents may lead to outcomes that fall short of AI4SG’s aim for positive and long-lasting change in social domains (Perera, 2024; Tudor et al., 2024). We highlight three prominent aspects below.
First of all, funding documents that motivated AI’s use with large-scale social problems without instantiating concrete contextual benefits (Section 4.1.1) could lead to premature emphasis on scalability, risking putting such a strong emphasis on creating something that works everywhere that it ends up not working particularly well anywhere. As Hanna and Park note, “scale thinking” necessitates homogenizing users into predefined categories, disregarding the diversity of local contexts (Hanna and Park, 2020). The urgency to scale and short-term funding could incentivize rushed implementations that are poorly suited to diverse contexts (Heeks, 2003, 2002; Dada, 2006). As Burrell and Toyama (2009) put it, “in an effort to solve problems quickly on large scale, a concern for local values, preferences, and opinions may become secondary,” cautioning against favoring only large-scale, generalizable studies at the expense of in-depth, context-rich work (and vice versa). We do not argue that scalable and generalizable approaches should not be pursued ever, but we suggest that they should be pursued inductively. That is, AI4SG initiatives should focus early efforts on solving specific problems of specific communities effectively and sustainably, and then look for commonalities that could inform the development of generalizable technical capabilities and infrastructures.
Secondly, funding rhetoric that led with assumptions of AI’s deterministic benefits without acknowledging its risks (Section 4.1.1) could reinforce the wider hype and “enchanted determinism” surrounding AI (Campolo and Crawford, 2020). This narrative fosters unrealistic expectations (Lin et al., 2024; Taylor et al., 2013), promotes techno-solutionism, and diverts attention from underlying social issues (Walsham, 2017; Karusala and Kumar, 2017). Furthermore, when funding documents framed AI as a force of good for the Global South or nonprofit sectors deterministically (Section 4.1.1), it risks promoting top-down approaches that cast these groups as lacking and in need of catching up (Erete et al., 2023; Valencia, 1997; Dearden and Tucker, 2016). This framing ignores the unique needs (and assets) of public sectors and the Global South (Hanna and Park, 2020). While corporate donations of computing resources and support (Section 4.4.1) can expand access, they may also reinforce Big Tech’s hegemonic narrative (Collins, 2000; Lukacz, 2024; Nost and Colven, 2022; Seger et al., 2023), addressing surface-level gaps rather than deeper structural issues (Ho et al., 2009b; Burrell and Toyama, 2009; Dearden and Tucker, 2016). These initiatives risk reproducing patterns of saviorism, positioning the Global South as a testing ground for foreign-owned systems while extracting valuable data (Damilare Dosunmu, 2025). These projects also overlook the uneven geography of the AI supply chain; the Global South is already part of AI, but primarily through the underpaid labor that makes AI systems work (Posada, 2022).
Last but not least, funding programs that targeted applicants with existing AI capacities without scaffolding for community engagement (Section 4.3.1) effectively shifted leadership power toward organizations or individuals who are technically skilled but likely lack a deep understanding of the social problem space. This emphasis also elevates AI knowledge production over community knowledge rooted in social contexts. This is fundamentally about epistemic justice—ensuring fairness in who is recognized as a “knower” and what forms of knowledge are legitimized (Fricker, 2007). Epistemic justice matters because when funding agendas prioritize technical capacity, they risk producing techno-centric solutions that lack community relevance, buy-in, and sustainability (Dodson et al., 2012).
5.2. Reframing AI4SG Funding: Recommendations for Funding Call Designs
Building on the more balanced approaches and existing gaps we identified in our dataset, we offer the following recommendations for funding call designs, putting them in conversation with existing literature on “tech for good” projects. We structure these recommendations around the four key components of funding calls, to serve as a template to inform the design of future funding calls.
5.2.1. Background: situating AI4SG in well-scoped contexts
Funding calls should orient towards projects that address well-scoped and contextually relevant problems. This doesn’t mean that funders should pre-define the problems, but they can support applicants by offering examples of relevant AI use cases or existing literature review on the social domain (Section 4.1.2). It may also involve requiring the active participation of communities during the problem formulation and solution brainstorming stage, where applicants and communities collaboratively investigate which technologies are appropriate for their specific problem. In a systematic review of community-collaborative approaches (CCA), Cooper et al. highlight that many projects involve communities only during limited stages (e.g., delivery of activities), rather than throughout the full research process—including problem framing and data analysis (Cooper et al., 2022). Yet, as Bratteteig and Wagner argue, engaging communities in deciding which problems to address is key, not only for inclusion but for redistributing power and ensuring they have real influence over project direction and outcomes (Bratteteig and Wagner, 2016). This level of participation reflects a deeper form of agency than simply evaluating or implementing pre-defined solutions (Delgado et al., 2023).
5.2.2. Expected Outcomes: defining success with intended communities
Successful outcomes in AI4SG projects should be defined collaboratively with the communities they aim to serve. Community perspectives on success often differ from funders’ or researchers’ goals, which tend to emphasize metrics like model accuracy or academic output (Lin et al., 2024; Erete et al., 2023)Lin et al. (2024) found that community organizations in AI4SG partnerships value learning about AI innovations, technical capacity building, and visibility of the projects even when real-world deployment isn’t achieved (Lin et al., 2024). Relatedly, Kumar and Dell (2018) report that development organizations prioritize accessible communication over formal research publications. Aligning success metrics early on helps manage expectations and reduces the risk of disappointment as projects progress (Taylor et al., 2013). While stakeholder goals may still sometimes conflict (Ho et al., 2009a), projects should aim for mutual benefit—for instance, through fair compensation for communities’ time and expertise or by supporting informal learning opportunities (Carroll and Rosson, 2013). This would also require attention to non-traditional (and often not quantifiable) metrics that capture relationship quality (Cooper et al., 2022; Erete et al., 2023). For example, funders could request letters of support and reports of experiences from communities as part of the project’s evaluation, alongside the quantitative evaluation of AI systems.
When real-world deployment is valued by the project teams and intended communities, it would be important to provide funding support and maintenance resources post-deployment—something notably absent from most funding programs in our dataset. In categorizing ICTD projects’ outcomes, Heeks calls attention to the high rate of projects that might have initially succeeded but break down over time (Heeks, 2002). As Jackson argues, “repair and maintenance” is a critical part of technology design, yet it pervasively falls out of view when the dominant discourse of technology focuses on innovations as a production, and not a process (Jackson, 2014). The maintenance work of a technology system is not only necessary for sustaining the system that will inevitably keep breaking (Jackson, 2014), it also builds relationships among workers who collaborate on the practical work (Irani and Silberman, 2014; D’Ignazio, 2024). Sustainable maintenance can involve building local capacity and intentionally transferring technology ownership to communities. For example, the research team behind SignSupport, a mobile app co-developed with the Deaf Community of Cape Town (DCCT) to support pharmacy interactions using South African Sign Language, trained community members to manage a computer lab and take on system administration and monitoring roles, while researchers stepped back to allow community leadership to grow without interference (Dearden and Tucker, 2016).
5.2.3. Eligibility and Team Composition: requiring community co-leadership and engagement beyond consultation
Funding calls should be structured to include intended community members as co-leaders or co-applicants, ensuring they have the power to shape project direction from the outset—and importantly, that they receive funding directly. As Bratteteig and Wagner point out, the “skill composition” of a design team significantly shapes how problems and solutions are defined (Bratteteig and Wagner, 2016). Including community members in core project roles enables them to define the problem space and success criteria, and ensures their labor and expertise are properly acknowledged and compensated (Pine et al., 2020). This approach also helps redistribute power within AI4SG collaborations, supporting more equitable relationships between communities and research institutions (Erete et al., 2023). Crucially, when engaging communities throughout the project, it is important to recognize and compensate for the various forms of labor they contribute—from sharing lived experience and domain expertise to facilitating connections within their communities. Future research could work towards quantifying both visible and invisible labor (see (Toxtli et al., 2021) and (Li et al., 2022)), while also acknowledging that some forms of labor in community engagement, such as emotional labor, may be difficult or impossible to measure.
5.2.4. Funders’ Support: valuing community engagement as a necessary part of AI4SG work
Funders can better support community engagement in AI4SG by recognizing that relationship-building is a foundational phase of the work, which often occurs before a clear project direction is even established (Carroll and Rosson, 2013; Cooper et al., 2022). This phase is essential to community-collaborative approaches (Cooper et al., 2022), and it requires project teams’ routine presence that takes time and emotional labor (Dearden and Tucker, 2016). Long-term relationships allow shared agendas to emerge, with community priorities guiding technology development from the start (Dearden and Tucker, 2016; Carroll and Rosson, 2013). One concrete step funders can take is to extend the typical AI4SG funding timeline (in our corpus: with a median of one and a half years) to accommodate this early-stage work. Relationship-building alone can take months (Erete et al., 2023) or even years (Carroll and Rosson, 2013; Dearden and Tucker, 2016), and should be supported as a formal project phase, even prior to defining problems or solutions. One key challenge is determining appropriate timelines and support—whether financial or capacity-building—for this work. Future research in CCA could analyze researchers’ reflections (e.g., (Erete et al., 2023; Dearden and Tucker, 2016; Carroll and Rosson, 2013)) to better understand the duration and nature of relationship-building across projects.
Furthermore, funders should also treat community engagement as a critical skill requiring training and support—on par with technical expertise. Many AI practitioners are trained in Computer Science departments that don’t routinely provide training in community engagement. As Okolo and Lin observe, AI4SG practitioners often use a hodgepodge of ad-hoc approaches to engage with end users (Okolo and Lin, 2024). Funders should recognize this knowledge gap and provide clear guidance for project teams to properly identify stakeholders or engage with end beneficiaries. There are multiple places where this guidance could be provided—as part of the funding call itself or additional training support that many funders already provide. Funders could consider citing existing guides in community engagement from the HCI community, such as the Equity-Centered Community Design Field Guide (Creative Reaction Lab, 2024), the Atlanta Community Engagement Playbook (Taranji Alvarado et al., 2024), and “readiness for partnership” checks (Pine et al., 2020). Other design toolkits like the Envisioning Cards (Friedman et al., 2024) might also be useful for AI4SG project teams to identify and engage with stakeholders. Since AI introduces additional hurdles to community-based designs, including requiring intensive data work (Miceli and Posada, 2022; Passi and Sengers, 2020; Sambasivan et al., 2021) and transparency issues (Okolo and Lin, 2024), HCI researchers could support funders by tailoring existing toolkits to the specific contexts of AI4SG, building on emerging Participatory AI4SG frameworks like the PACT framework (Bondi et al., 2021).
5.3. Calling for Meaningful Engagement: Recommendations for the Funding Design Process
Beyond the design of individual funding calls, we offer recommendations for funders on the design process of funding programs. The design process of grantmaking is influenced by funders’ connections to other stakeholders (Saha et al., 2022a). From our dataset, we observed that private tech companies were not only the providers of AI products, resources, and training support, but also AI4SG funders themselves. Through investment in AI4SG, private tech companies could not only expand their market influence (Henriksen and Richey, 2022; Lukacz, 2024; Nost and Colven, 2022) but also influence other funders’ beliefs about AI’s potential and their funding priorities. At the same time, the influence of communities on funding documents remains scant (c.f.  (Saha et al., 2022a)). To reorient AI4SG towards a more balanced approach, we see the need for funders to engage with intended communities and the HCI community.
5.3.1. Engaging with intended communities
Engaging with communities in the design of funding calls would enable funders to align their agendas with the needs and visions of those communities early in the decision-making process, before projects even begin (Saha et al., 2022a; Balestrini et al., 2014). For funding programs that target projects in a region that is far away from the funding organizations’ locations, it would be important to have regional partners with a more solid contextual understanding, such as regional research hubs and organizations. They can serve as “intermediaries” between funding organizations and intended end beneficiaries that funding organizations don’t have direct access to (Dell and Kumar, 2016; Saha et al., 2022a).
However, there are several challenges that funders might face in engaging with communities. Gardner et al. point out that staff in funding organizations don’t necessarily have internal expertise in ethical AI development to craft and incorporate responsible AI principles in public AI funding (Gardner et al., 2022). Relatedly, through interviews with decision-makers from grant foundations for ICTD projects, Saha et al. found that while funders often have a desire to connect with communities, they struggle to incorporate community voices into the grantmaking process, due to barriers such as geographical distances, language barriers, and a lack of institutional incentives (Saha et al., 2022a). Engaging with the HCI community could help address some of these challenges in incorporating community voices into the grantmaking process.
5.3.2. Engaging with the HCI community
Following up on Saha et al.’s call on HCI researchers and practitioners to serve as “mediators” between decision-makers and communities’ voices (Saha et al., 2022b), we see opportunities in the AI4SG space where HCI researchers and practitioners can mediate between communities, funders, and other stakeholders like technologists and funders’ regional partners. Specifically, the HCI community can support funders in identifying key stakeholders to engage with. HCI practitioners could also help funders design training materials on community engagement for project teams and provide community engagement workshops throughout the project stages. While there is a rich body of resources for community-collaborative approaches (CCA) (Cooper et al., 2022), there exist gaps for the HCI community to bridge between frameworks and practices, such as implementing formal evaluations of participatory design projects (Bossen et al., 2016; Nguyen, 2022).
Moreover, the HCI community can provide tools and facilitations to foster reflexivity in the grantmaking process, such as reflection workshops that bring funders, regional partners, representatives of intended beneficiaries, and targeted applicants together to learn from past funding program experiences. Wellcome Trust’s design process of Mental Health Data Prizes (C8 and C9) was a good example of benefits of feedback workshops (Section 4.4.2). HCI researchers and practitioners can provide consultations on reflection workshops or deliberation tools when considering strategies around AI (e.g., (Kawakami et al., 2024b)).
Last but not least, the HCI community can offer counter-narratives to the more techno-centric conceptualizations of AI in funding calls, assisting funders to consider alternatives that emphasize diverse ways of knowing and sustainable development of AI within planetary boundaries. Some emerging critical reassessments of current AI design practices and public discourse around AI include work on AI abandonment (Johnson et al., 2024), post-capitalist AI systems based on feminist values (Varon et al., 2024), Indigenous thoughts (Ofosu-Asare, 2024; Lewis et al., 2024), post-colonial computing (Sultana and Ahmed, 2019), and post-growth philosophies (Sharma et al., 2023; Kuijer and Laschke, 2024).
5.4. Limitations and Future Work
The list of funding documents in our dataset is neither exhaustive nor statistically representative. Our analysis is restricted to English-language materials and funders located in English-speaking regions, which excludes funding programs described in other languages. Additionally, funding programs that do not use the keywords in our study might have been overlooked. Future work could expand the analysis to AI4SG funding documents published in languages other than English. Furthermore, while calls for proposals and grant announcements reveal funders’ intentions, they do not necessarily reflect what projects actually get funded. Building on this work, future research could examine the characteristics of funded projects. This line of work could reveal alignment or gaps between vision and practice, including who receives funding (whether AI-capable organizations or mission-driven groups). Future work could also explicitly examine project outcomes to categorize commonalities of success and failures specific to AI4SG (similar to Dodson et al. (2012)’s analysis of ICTD projects). Since publicly available written materials may be biased toward success stories, focusing on one or a few case studies—with interviews with project teams and analysis of project-related documents—could yield insightful analysis. Finally, we emphasize that the funding documents in our corpus fall along a spectrum from more techno-centric to more balanced approaches, and we caution readers against viewing this spectrum as a dichotomy. Moreover, future work could further explore what it may look like to over-emphasize contextual understanding and under-emphasize technical capacity.
6. Conclusion
Artificial Intelligence for Social Good (AI4SG) emerges as a growing body of research and practice that explores the potential of AI and data-driven technologies to tackle social problems like public health and humanitarian aid (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). However, a growing body of evidence suggests that AI4SG initiatives do not necessarily yield the intended positive outcomes in real-world contexts (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). While prior work has focused on the actions and roles that project teams—typically AI developers and their nonprofit partners–play in project outcomes, an emerging body of work points to the significant influence of funders in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Gardner et al., 2022). While prior work has highlighted funders’ significant influence, we still lack a direct understanding of how funders frame AI’s role in addressing social good challenges and how they envision AI’s impact could be achieved.
In this study, we fill this gap and “study up” AI4SG funding documents by examining them through the lens of existing literature on “tech for good” projects. Our analysis reveals dissonances between AI4SG funders’ intentions for positive, sustainable social impact and the techno-centric approaches that some funding programs promoted. At the same time, we identified several more balanced approaches as well as existing gaps in funding calls. Building on our analysis, we offer recommendations for funders in funding call designs, while emphasizing the need for community engagement in the funding call design process. Finally, we call on the HCI community and community-collaborative approaches practitioners and researchers to mediate between AI4SG stakeholders to incorporate meaningful engagement with communities in the early stages of AI4SG projects—when funding programs are being designed.
Appendix A Funding documents
See Table LABEL:tab:funding_doc_mr next page.
Table 1. Funding Documents Related to AI and Social Good
ID Title Year Funding Amount Funding Duration Funder(s) Funder Types Funder HQ Eligible Regions Social Domains Eligible Applicants Ref.
C1 Artificial Intelligence (AI) for Societal Good Science Foundation Ireland Future Innovator Prize 2019 multi-phased (Concept: €20K; Seed: €200K; Prize: €2M) multi-phased (Concept: 3 months; Seed: 9 months; Prize: 12 months) Science Foundation Ireland Government Agency Ireland Ireland UN SDGs “Academic Staff or Contract Researcher; Holds a PhD (or equivalent)” (Science Foundation Ireland, 2019)
C2 UNICEF Innovation Fund Call for Data Science & A.I 2019 \$100,000 per project not specified UNICEF Non-profit United States Global broadly “benefit humanity” “private company in a UNICEF programme country” (UNICEF, 2019)
C3 Google AI Impact Challenge (Continued as AI for the Global Goals) 2018 \$25 million in total (for 20 organizations) 1–3 years Google Private Tech Company United States Global broadly “social and environmental challenges” “Organizations that are experienced in AI to those with an idea for how they could put their data to better use” (Google, 2018)
C4 A Grand Challenges Request for Proposals: Catalyzing Equitable Artificial Intelligence (AI) Use 2023 \$100,000 per project (\$3,000,000 total) up to 3 months Bill & Melinda Gates Foundation Philanthropy Foundation United States Low- and Medium-Income Countries (LMICs) Decision-making in LMICs “investigators in LMICs” (The Bill & Melinda Gates Foundation, 2023)
C5 Innovative Data and Modeling Approaches to Measure Women’s Health 2025 Up to \$150,000 per project 18–24 months Bill & Melinda Gates Foundation Philanthropy Foundation United States Global (preferring Sub-Saharan Africa and South Asia) Women’s Health “nonprofit organizations, for-profit companies, international organizations, government agencies and academic institutions” (Bill & Melinda Gates Foundation, 2025)
C6 Data for Social Sustainability call for proposals 2021 Up to \$80,000 (avg) + AWS credits (up to \$20,000 avg) 1 year Amazon Private Tech Company United States Global human rights protections “full-time faculty member or permanent NGO researcher” (Amazon, 2021)
C7 Mozilla Technology Fund: AI and Environmental Justice 2024 \$50,000 per project (\$300,000 total) 1 year Mozilla Foundation Non-profit United States Global environmental justice “all applicants regardless of geographic location or institutional affiliation” (Mozilla Foundation, 2023)
C8 Wellcome Mental Health Data Prize: UK and South Africa 2022 multi-phased (Discovery: £40,000; Prototyping: £100,000; Sustainability: £500,000) multi-phased (each phase 6 months) Wellcome Trust Philanthropy Foundation United Kingdom United Kingdom and South Africa anxiety and depression in young people “higher education, research institutes, non-academic healthcare orgs, not-for-profits” (Wellcome Trust, 2022)
C9 Wellcome Mental Health Data Prize: Africa 2024 £200,000 per team (10 teams total) 1 year Wellcome Trust Philanthropy Foundation United Kingdom Africa anxiety, depression, and psychosis “lead applicants based in Africa (academic, non-profits on mental health, government agencies, research institutions)” (APHRC, 2024)
C10 Request for Applications (RFA) for AI innovations in MSRH 2022 \$10,000–\$15,000 (students/startups); \$30,000–\$40,000 (established orgs) 1.5 years IDRC and Sida Government Agency Canada and Sweden Sub-Saharan Africa maternal, sexual, and reproductive health/rights “students and start-up organisations; established organisations” (HASH, 2022)
C11 Call for proposals: AI for climate action Innovation Research projects 2022 \$50,000–\$60,000 1.5 years IDRC and Sida Government Agency Canada and Sweden Africa climate change “researchers from accredited universities; public/private research institutions” (RUFORUM and WASCAL, 2022)
C12 Innovate Africa Challenge Artificial Intelligence (AI) for Climate Action 2024 \$50,000 multi-phased (bootcamp; incubation; implementation) BMZ Government Agency Germany Africa climate change “company from any of the 39 Smart Africa Alliance countries” (Smart Africa, 2024)
C13 Call for Proposals: Climate Change AI Innovation Grants 2024 up to \$150,000 per proposal (up to \$1.4M total) 1 year Climate Change AI (QCF, Google DeepMind, Global Methane Hub) Non-profit United States Global climate change “PI affiliated with accredited university in one of 38 OECD member countries” (Climate Change AI, 2024)
C14 Request for Proposals: Natural Language Processing (NLP) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Google.org) Non-profit United States Africa and Latin America NLP data “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024a)
C15 Request for Proposals: Antimicrobial Resistance (AMR) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Wellcome Trust) Non-profit United Kingdom LMICs in Africa, Latin America, South & Southeast Asia public health “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024b)
C16 Public-Private Partnerships to Improve Population Health Using AI/ML 2025 Up to \$612,500 per project (Planning+Implementation) 18 months NIH Government Agency United States United States public health “Public-Private dyad or triad applicant teams” (AIM-AHEAD, 2025)
C17 Big Data Science: Mechanisms of HIV Activity & Interaction with Substance Use Disorder 2023 \$2,000,000 Up to 5 years NIH Government Agency United States Global public health Organizations (National Institutes of Health (2025), NIH)
C18 DS-I Africa Research Hubs 2020 \$32,500,000 Up to 5 years NIH Government Agency United States Africa public health “Non-domestic (non-U.S.) Entities (Foreign Institutions)” (National Institutes of Health (2020), NIH)
C19 Smart Health & Biomedical Research in the Era of AI & Advanced Data Science (SCH) 2025 Up to \$20,000,000 per year Up to 4 years NSF Government Agency United States United States public health “Institutions of Higher Education; Non-profit, non-academic organizations” (U.S. National Science Foundation, 2025)
A1 “Patrick J. McGovern Foundation commits \$40M to advancing AI and data for good” 2021 \$40,000,000 not specified Patrick J. McGovern Foundation Philanthropy Foundation United States unspecified broadly nonprofits “Nonprofits and NGOs; orgs on data dignity, data stewardship, AI ethics” (Foundation, 2021)
A2 “Artificial Intelligence to Accelerate Inclusion Challenge” 2024 not specified not specified Mastercard Center for Inclusive Growth; data.org Non-profit United States Global economic empowerment “the diverse range of communities often left out of AI for social impact” (Mastercard, 2024)
A3 “DataKind to help more organizations use data for social change with new \$1.1 million investment from Knight Foundation” 2015 \$1,100,000 3 years John S. and James L. Knight Foundation Philanthropy Foundation United States Global broadly nonprofits DataKind (Knight Foundation, 2015)
A4 “AI for local news: advancing business sustainability in newsrooms” 2021 \$3,000,000 2 years John S. and James L. Knight Foundation Philanthropy Foundation United States United States local journalism “local news organizations” (Knight Foundation, 2021)
A5 “The Rockefeller Foundation Invests in Satellite Data & AI to Accelerate Economic Development and Climate Resilience in Africa” 2022 \$5,500,000 3 years The Rockefeller Foundation Philanthropy Foundation United States United States development e-GUIDE and Atlas AI (The Rockefeller Foundation, 2022b)
A6 “The Rockefeller Foundation Commits \$1M to Expand BlueConduit’s AI-Based Solution to America’s Lead Pipe Challenge” 2022 \$1,000,000 5 years The Rockefeller Foundation Philanthropy Foundation United States United States infrastructure BlueConduit (The Rockefeller Foundation, 2022a)
A7 “Combining forces for a new phase of AI for development: Africa and beyond” 2023 CAD130,000,000 5 years IDRC; UK FCDO; Bill & Melinda Gates Foundation; USAID Government Agency, Philanthropy Foundation Canada; United Kingdom; United States Africa development unspecified (IDRC, 2023)
A8 “Artificial Intelligence for Development (AI4D)” 2020 CAD 20,000,000 5 years IDRC and Sida Government Agency Canada and Sweden Africa development unspecified (IDRC, 2020)
A9 “UK unites with global partners to accelerate development using AI” 2023 £38,000,000 5 years the UK government Government Agency United Kingdom Africa development unspecified (UK Government, 2023)
A10 “data.org Announces Generative AI Skills Challenge Awardees” 2023 \$250,000 per organization (5 orgs) not specified data.org; Microsoft Private Tech Company United States Global economic empowerment Organizations broadly (data.org, 2023)
A11 “Inclusive Growth and Recovery Challenge” 2021 \$10,000,000 not specified data.org; Mastercard Center for Inclusive Growth; The Rockefeller Foundation; The Paul Ramsay Foundation Non-profit United States; Australia unspecified development unspecified (data.org, 2021)
A12 “Announcing AI for Earth: Microsoft’s new program to put AI to work for the future of our planet” 2017 \$2,000,000 1 year Microsoft Private Tech Company United States Global sustainability “researchers and organizations” (Smith, 2017b)
A13 “AI for Earth can be a game-changer for our planet” 2017 \$50,000,000 5 years Microsoft Private Tech Company United States Global sustainability “universities, nongovernmental organizations and others” (Smith, 2017a)
A14 “Using AI to help save lives” 2018 \$40,000,000 5 years Microsoft Private Tech Company United States Global humanitarian crisis “nongovernmental organizations (NGOs) and humanitarian organizations” (Smith, 2018)
A15 “Google.org Impact Challenge: Tech for Social Good” 2023 €16,000,000 up to 6 months Google Private Tech Company United States Europe sustainability, economic opportunity, cybersecurity “European organisations with projects focused on sustainability, economic opportunity or cyber security” (Google, 2023)
A16 “Google.org Accelerator: Generative AI open call” 2025 \$30,000,000 6 months Google.org Private Tech Company United States Global broadly nonprofits social impact organizations (Google.org, 2025)
Appendix B Coding Process and Affinity Diagram
Figure 6. The first two authors independently conducted an open-ended inductive qualitative analysis of three randomly chosen funding documents. Through iterative discussions and collective affinity diagramming on an online whiteboard, the two authors identified some initial themes, which helped refine our research questions and coding strategies.
Figure 7. After an open-ended inductive coding process (Figure 6), we narrowed the scope of analysis to four main components of funding documents: 1) Background, 2) Expected Outcomes, 3) Eligibility Criteria and Team Composition, and 4) Funders’ Support. We characterized high-level themes into a spectrum between a more techno-centric approach or a more balanced approach through collective affinity diagramming.